trains a and b are traveling in the same direction on parallel tracks. train a is traveling at 80 miles per hour and train b is traveling at 100 miles per hour. train a passes a station at 3:10 pm. If train B passes the same station at 3:22 pm at what time will train b catch up to a.

Answer by Alan3354(6097)

trains a and b are traveling in the same direction on parallel tracks. train a is traveling at 80 miles per hour and train b is traveling at 100 miles per hour. train a passes a station at 3:10 pm. If train B passes the same station at 3:22 pm at what time will train b catch up to a.
----------------------
In the 12 mins from 310 to 322, train a goes 16 miles (80 * 1/5 hour)
Train B gains on train A at 20 mph (100 - 80)
16/20 = 4/5 hours = 48 minutes
322 + 48 = 4:10 PM